# Feature Identification via the Empirical NTK

### v2

@@ Abstract
We provide evidence that eigenanalysis of the empirical neural tangent kernel (eNTK) can surface the features used by trained neural networks. Across two standard toy models for mechanistic interpretability, Toy Models of Superposition (TMS) and a 1-layer MLP trained on modular addition, we find that the eNTK exhibits sharp spectral cliffs whose top eigenspaces align with ground-truth features. In TMS, the eNTK recovers the ground-truth features in both the sparse (high superposition) and dense regimes. In modular arithmetic, the eNTK can be used to recover Fourier feature families. Moreover, we provide evidence that a layerwise eNTK localizes features to specific layers and that the evolution of the eNTK spectrum can be used to diagnose the grokking phase transition. These results suggest that eNTK analysis may provide a practical handle for feature discovery and for detecting phase changes in small models.
@@ 1 Introduction
Mechanistic interpretability seeks to understand how neural networks process information at inference time (Olah et al. (2020)). A major open question is to understand how neural networks represent learned features. Operationally, suppose that an input to a model can either exhibit a particular human-interpretable property (e.g. the color red, a dog snout) or not. Neural networks often seem to learn to detect such properties. For example, they can be probed to classify the presence or absence of such properties after pretraining with higher-than-chance accuracy (Alain and Bengio (2016)). We would like to understand how they do so internally, by finding functions of a model's activations and weights that detect the presence of the property with high sensitivity and specificity.
It's a priori unclear how to ``guess" a good candidate answer. One strategy is to look to theories of deep learning. If there existed a hypothetical complete theory that allowed us to ``reverse-compile" any model into human-interpretable code, it would (by definition) solve this problem by telling us which effective features a model uses and how they interact to produce the input-output map (Olah (2022)). Instead of a complete theory however, at present we have several nascent theories that each explain some modest aspects of what neural networks may be doing, in some corners of the parameter space.  [^1] See e.g. the examples on page 13 of Sharkey et al. (2025).
Among such theories in 2025, the neural tangent kernel (NTK) and its cousins (Jacot et al. (2018), Roberts et al. (2022)) are unique in giving us a closed-form formula for the function $f$ learned by a neural network in a corner of the parameter space: namely
$f_{i}(x)=\sum_{\alpha_{1},\alpha_{2}\in\mathcal{D}}K_{ij}(x,x_{\alpha_{1}})K^{-1}_{jk}(x_{\alpha_{1}},x_{\alpha_{2}})\,y_{k,\alpha_{2}}\,$, (1)
where the index $i$ runs over output neurons (classes), the $\alpha'$s run over points in the training set $\mathcal{D}$, and the data-space kernel
@@
evaluated at initialization, with the index $\mu$ running over model parameters, is the eponymous NTK. Eq. (1) is a provably good approximation to the function learned by a neural network under gradient descent with MSE loss in a ``lazy limit" where the drift in model parameters throughout training is parametrically small. See Appendix A for a review.
Realistic models are usually not in this regime. However, it's been conjectured and empirically verified in some situations (Fort et al. (2020), Atanasov et al. (2021), Ortiz-Jiménez et al. (2021), Mohamadi et al. (2023)) - although we don't yet have a general understanding of when the conjecture is and isn't true - that Eq. (1) with the NTK (2) evaluated at the end of training instead of at initialization,
@@
can at times be a good approximation to the function learned by a neural network beyond the lazy limit. This conjecture goes by the name of the empirical NTK hypothesis (eNTK hypothesis), with (3) being the empirical NTK (eNTK).
@@
In this paper, we will take this conjecture seriously and see if the eNTK can be used to generate new algorithms for mechanistic interpretability. In particular, we examine the hypothesis that eigenvectors of the eNTK may align with interesting features at the end of training. We find that this idea turns out to be true in toy models commonly studied by the mechanistic interpretability community. Our main contributions are that
@@
  • In ``Toy Models of Superposition" (Elhage et al. (2022)), the eNTK eigenspectrum is anisotropic with a qualitatively large drop at the $n$th ordered eigenvalue, for $n$ the number of ground-truth features. The leading $n$-dimensional eigenspace aligns nearly 1:1 with the ground-truth features at the end of training in both the sparse and dense regimes.
  • For a 1L MLP trained on modular arithmetic mod $p$ in the grokking regime, the eNTK eigenspectrum develops two cliffs. The first cliff, which is localized to the first layer of the model and visible already at initialization, has dimension $k=4\lfloor p/2\rfloor$ and a span that matches the Fourier features learned by the first layer of the model. A Laplacian (axis-smoothness) rotation inside this subspace yields basis vectors that exactly align with the Fourier features used by the model.
@@
  • The second cliff aligns with ``sum" and ``difference" Fourier features and appears precisely at the grokking phase transition. In this sense, eNTK spectral analysis provides a practical diagnostic that localizes where features live by layer and detects phase transitions.
@@
This paper is organized as follows. In Section 2, we explain how we use the eNTK to find features in trained models. In Sections 3 and 4, we use the eNTK to find features in Toy Models of Superposition and a 1L MLP trained on modular arithmetic, respectively. We conclude with a discussion of future directions in Section 5.
@@
Related work. A few earlier works explored whether NTK eigenvectors align with interesting directions in trained models. Loo et al. (2022) and Tsilivis and Kempe (2022) visualized kernel eigenvectors as candidate features in small image nets (in the context of using the NTK to study adversarial training), and are the most closely related works. Atanasov et al. (2021) and Baratin et al. (2021) argued that NTK tangent features rotate towards directions predictive of the outputs during training.
@@
Besides the eNTK, other objects involving gradients/curvature have appeared in the mechanistic interpretability literature. As mentioned above, the eNTK (3) is constructed by contracting a pair of Jacobians $J(z)=\nabla_{W}f(z)$ along the parameter direction, forming a kernel in data space (shorthand: $K(z,z^{\prime})=J(z)J(z^{\prime})^{T}=\nabla_{W}f(z)\nabla_{W}f(z^{\prime})^{T}$). The Hessian, the second derivative of the loss function wrt parameters $H=\nabla^{2}_{W}\mathcal{L}$, can be approximated by a pair of Jacobians contracted the other way ($H\approx\sum_{i}J(x_{i})^{T}\,\nabla^{2}_{f}\mathcal{L}(f(x_{i}),y_{i})\,J(x_{i})$), and appears frequently in recent work on singular learning theory (e.g Bushnaq et al. (2024b)). The Hessian induces a ``influence function" kernel on data points $\mathcal{I}(z^{\prime},z)=-\nabla_{W}\mathcal{L}(z^{\prime})^{T}H^{-1}\nabla_{W}\mathcal{L}(z)$ (Koh and Liang (2017)) that was recently used to interpret LLMs at scale (Grosse et al. (2023)). The Jacobian also played a role in the ``Local Interaction Basis" algorithm (Bushnaq et al. (2024a)) which proposed to find interpretable directions in activation space in part by aligning would-be directions with singular directions of the intra-layer Jacobian. Somewhat further afield, attribution patching (Nanda (2023)) approximates activation patching (an interpretability technique where one swaps clean activations into a corrupted forward pass to test causal effect) by estimating the effect using the activation's local gradient.
Non-gradient-based kernels have also appeared in the interpretability literature. In SAE work (e.g. Bricken et al. (2023), Templeton et al. (2024)) it's common to compute a kernel-like measure of similarity between SAE features (e.g. cosine similarity of their neuron supports), then use nearest-neighbors/UMAP to visualize the feature manifold. More recently, motivated by singular learning theory, Adam et al. (2025) used a data-space ``loss kernel" defined as the covariance of per-sample losses under low-loss perturbations of a model's weights to cluster inputs/map dataset geometry.
@@
The idea of approximating a neural network by keeping only kernel modes up to an emergent spectral scale can be viewed as a naive application of renormalization ideas to mechanistic interpretability (Greenspan (2025)), with the spectral scale setting a hard UV cutoff.
2 Method: NTK pipeline for feature-finding
@@
The main idea that we pursue in this paper is to treat the eNTK as fundamental and interpret its top eigenvectors. One motivation for doing so is that empirically, the spectrum of the eNTK is often very anisotropic. (We'll see many examples of this below.) When true, it means that if the linearized formula (1) is a good approximation to the function learned by a neural network at the end of training, the same formula where we replace the kernel with its low-rank spectral truncation should in turn be a good approximation, and the top eigenvectors of the kernel should contain most of the information about how the trained model computes its outputs.
Below, for several models, we will eigendecompose the empirical NTK and its variants defined in section 2.1. We define a cliff to be a (qualitatively) large drop in the ordered eigenvalues and a NTK feature subspace to be the span of eigenvectors corresponding to the eigenvalues in a cliff.
@@
To compute if a length-$N$ eNTK eigenvector for $N$ the dataset size aligns with a ground-truth feature, we can take the absolute value of its cosine similarity with a length-$N$ ``feature vector" whose entries consist of that feature's activation on each data point. This gives a scalar for each (eigenvector, feature) pair. To visualize how the eigenvectors across a NTK feature subspace align with a selection of features, we will render heatmaps that show this correlation across all pairs.
2.1 Variants of the eNTK used in this paper
@@
The eNTK evaluated over all points in a dataset has shape $(N,N,C,C)$ for $N$ the size of the dataset, and $C$ the number of output neurons (classes). To do an eigendecomposition, we must reduce it to two dimensions. Below we'll use two collapses of $K$:
@@
Another variant that we'll use is the layerwise NTK where rather than sum over all model parameters in (3), we sum only over those parameters belonging to a particular layer, before applying either collapse above. This will prove useful for attributing features to a layer.
3 Results for ``Toy Models of Superposition"
In our first experiment, we study the alignment of eNTK eigenvectors with ground-truth features in Toy Models of Superposition (TMS) (Elhage et al. (2022)).
@@ 3.1 Review of the setup
the dataset consists of NN length-$n$ vectors whose $i$th entry is set to 0 with probability $S$, and otherwise randomly drawn from $U[0,1]$  [^2] By using a fixed-size dataset, we’re actually using the setup of Henighan et al. (2023). ; and the loss function is MSE loss modified by a per-feature importance hyperparameter $I_{i}$,
@@
The claim to fame of this setup is that it's a maximally simple toy model of superposition: namely, the idea that if models store features as linear combinations of within-layer neurons, then a model might learn to represent more features than neurons available to it by storing them non-orthogonally in the neural activations. In this experiment, the $n$ input/output directions are the analogs of ground-truth features, and one would like to see if the autoencoder is able to reconstruct them. Upon training this class of model, one finds that when the features are dense (small $S$), the model learns to reconstruct the top $m$ features only, but as one dials up the sparsity, the model is able to recover more and more features, at the cost of encoding them non-orthogonally.
@@
Figure 1: Eigenvalue spectrum of the flattened eNTK (top row) and column-normalized, importance-rescaled flattened eNTK eigenvector-feature indicator heatmaps for TMS models with $n=50$ ground-truth features, importance $I_{i}=0.8^{i}$, (hidden-layer size, sparsity) $m=10,S=0.3$ (left column); $m=10,S=0.9$ (middle column) or $m=40,S=0.9$ (right column); and importance rescaling parameters $\beta=1$ (middle row) and $\beta=0.3$ (bottom row) used as described in the main text. The center column is an example of a TMS model in the sparse regime (high superposition), while the left and right columns are two different ways to push the model to the dense regime (low superposition).
@@ 3.2 Results
Spectral structure and cliffs. Across TMS settings (dense vs. sparse regime and different widths $m$), the flattened eNTK exhibits a pronounced spectral cliff structure separating a small set of leading modes from a long tail (Fig. 1, row 1). In all cases that we tested, the leading eigenspace contains two cliffs with the second drop appearing at $k=50$, equal to the number of ground-truth features in the dataset. A first cliff differentiates the number of features that the model is able to reconstruct from those it cannot.  [^3] Quantitatively, in each example the number of entries along the diagonal of $W^{T}W$ exceeding the heuristic threshold of 0.75, for $W$ the trained weight matrix, equals the size of one of the two cliffs.
@@
Alignment with ground-truth features. As discussed in section 2, one way to visualize how well a length-$N$ eigenvector for $N$ the dataset size lines up with a ground-truth feature is to compute its inner product with a ``feature vector" whose entries consist of that feature's activation on each data point. Here, we choose instead to work with the flattened eNTK which is a shape $NC$ by $NC$ tensor. At the same time, the output classes $C$ correspond to ground-truth features in TMS. To construct heatmaps displaying how well the flattened eNTK eigenvectors align with ground-truth features in TMS, we therefore first build an expanded data matrix of shape $(NC,C)$ whose entries are zero unless the class indices match, then compute and display the inner product of the expanded data matrix with the top flattened eNTK eigenvectors.
@@
Before rendering the heatmaps, we also apply an importance rescaling to the eNTK, multiplying each Jacobian entry $df_{i}(x)/dW_{\mu}$ in (3) by $I^{\beta i/2}$. This directionally corrects for the fact that the derivation of Eq. (1) assumed a standard MSE loss (see Appendix A), but TMS uses an importance-weighted MSE loss, (5). The choice $\beta=1$ is the theoretically principled one that absorbs $I$ into the NTK in eq. (13), allowing the rest of the derivation in Appendix A to proceed with standard MSE, while $\beta=0$ corresponds to turning off the importance rescaling.
@@
With $\beta=1$, we find that the leading eigenvectors align with high-importance features (middle row of Fig. 1). We also sweep over other values of $\beta\in[0,1]$ and find that as $\beta$ decreases, lower-importance features become increasingly visible in the heatmaps. In particular, at the value $\beta=0.3$ (bottom row of Fig. 1), the leading eigendirections of the importance-rescaled eNTK appear to align nearly one-to-one with all of the ground-truth features of TMS in both the sparse and dense regimes. This experiment suggests that the raw kernel also contains information about less-important features, but down-weighs them in line with the objective.
@@
In the sparse (high superposition) regime, the map between top eNTK eigenvectors and ground-truth features appears to split into two lines sorted by importance (center and bottom middle plots in Fig. 1). It could be interesting to try to understand why this happens. This is likely possible given the simplicity of the kernel itself in TMS (see Appendix B for the closed-form expression of the kernel), but we will leave it for future work.
4 Results for modular arithmetic (grokking)

### v3

@@ Abstract
We provide evidence that eigenanalysis of the empirical neural tangent kernel (eNTK) can surface the features used by trained neural networks. Across three standard toy models for mechanistic interpretability, Toy Models of Superposition (TMS), a 1-layer MLP trained on modular addition and a 1-layer Transformer trained on modular addition, we find that top eigenspaces of the eNTK align with ground-truth features. In TMS, the eNTK recovers the ground-truth features in both the sparse (high superposition) and dense regimes. In modular arithmetic, the eNTK can be used to recover Fourier feature families. Moreover, we provide evidence that a layerwise eNTK localizes features to specific layers and that the evolution of the eNTK spectrum can be used to diagnose the grokking phase transition. These results suggest that eNTK analysis may provide a practical handle for feature discovery and for detecting phase changes in small models.
Machine Learning, ICML
@@ 1 Introduction
Mechanistic interpretability seeks to understand how neural networks process information at inference time (Olah et al., 2020). A major open question is to understand how neural networks represent learned features. Operationally, suppose that an input to a model can either exhibit a particular human-interpretable property (e.g. the color red, a dog snout) or not. Neural networks often seem to learn to detect such properties. For example, they can be probed to classify the presence or absence of such properties after pretraining with higher-than-chance accuracy (Alain & Bengio, 2016). We would like to understand how they do so internally, by finding functions of a model’s activations and weights that detect the presence of the property with high sensitivity and specificity.
It’s a priori unclear how to guess a good candidate answer. One strategy is to look to theories of deep learning. If there existed a hypothetical complete theory that allowed us to “reverse-compile” any model into human-interpretable code, it would (by definition) solve this problem by telling us which effective features a model uses and how they interact to produce the input-output map (Olah, 2022). Instead of a complete theory however, at present we have several nascent theories that each explain some modest aspects of what neural networks may be doing, in some corners of the parameter space.  [^1] See e.g. the examples on page 13 of (Sharkey et al., 2025).
Among such theories in 2025, the neural tangent kernel (NTK) and its cousins (Jacot et al., 2018; Roberts et al., 2022) are unique in giving us a closed-form formula for the function $f$ learned by a neural network in a corner of the parameter space: namely
$f_{i}(x)=\sum_{\alpha_{1},\alpha_{2}}K_{ij}(x,x_{\alpha_{1}})K^{-1}_{jk}(x_{\alpha_{1}},x_{\alpha_{2}})\,y_{k,\alpha_{2}}\,$, (1)
where the index $i$ runs over output neurons (classes), the $\alpha$’s run over points in the training set, and the data-space kernel
@@
evaluated at initialization, with the index $\mu$ running over model parameters, is the eponymous NTK. Eq. (1) is a provably good approximation to the function learned by a neural network under gradient descent with MSE loss in a “lazy limit” where the drift in model parameters throughout training is parametrically small. See Appendix A for a review.
Realistic models are usually not in this regime. However, it’s been conjectured and empirically verified in some situations (Fort et al., 2020; Atanasov et al., 2021; Ortiz-Jiménez et al., 2021; Mohamadi et al., 2023) - although we don’t yet know in general when the conjecture is and isn’t true - that Eq. (1) with the NTK (2) evaluated at the end of training instead of at initialization,
@@
can at times be a good approximation to the function learned by a neural network beyond the lazy limit. This conjecture goes by the name of the empirical NTK hypothesis, with (3) being the empirical NTK (eNTK).
@@
In this paper, we will take this conjecture seriously and see if the eNTK can be used to generate new algorithms for mechanistic interpretability. In particular, we examine the hypothesis that eigenvectors of the eNTK may align with interesting features at the end of training. We find that this turns out to be true across toy models commonly studied by the mechanistic interpretability community. Our main contributions are that
@@
  • In “Toy Models of Superposition” (Elhage et al., 2022), the eNTK eigenspectrum is anisotropic with a qualitatively large drop at the $n$th ordered eigenvalue, for $n$ the number of ground-truth features. The leading $n$-dimensional eigenspace aligns nearly 1:1 with the ground-truth features at the end of training in both the sparse and dense regimes.
@@
  • For a 1L MLP trained on modular arithmetic mod $p$ in the grokking regime, the eNTK eigenspectrum develops two cliffs. The first cliff, localized to the first layer of the model and visible already at initialization, has dimension $k=4\lfloor p/2\rfloor$ and a span that matches the Fourier features learned by the first layer of the model. The second cliff aligns with the “sum” and “difference” Fourier features learned by the second layer of the model and appears precisely at the grokking phase transition. In this sense, eNTK spectral analysis provides a practical diagnostic that localizes where features live by layer and detects phase transitions.
  • For a 1L Transformer trained on modular arithmetic mod $p$ in the grokking regime, across training runs, the leading eigendirections of the eNTK localized to the O/V layers of the attention block, the first (input) layer of the MLP block, the second (output) layer of the MLP block and the unembedding layer all align with Fourier features at the run-dependent “key frequencies” used by the model to implement the modular addition algorithm.
@@
This paper is organized as follows. In Section 2, we explain how we use the eNTK to find features in trained models. In Sections 3, 4 and 5, we use the eNTK to find features in Toy Models of Superposition, a 1L MLP trained on modular arithmetic and a 1L Transformer trained on modular arithmetic, respectively. We conclude with a discussion of future directions in Section 6.
@@
Related work. A few earlier works have explored whether NTK eigenvectors align with interesting directions in trained models. (Loo et al., 2022) and (Tsilivis & Kempe, 2022) visualized kernel eigenvectors as candidate features in small image nets (in the context of using the NTK to study adversarial training), and are the most closely related works. (Atanasov et al., 2021) and (Baratin et al., 2021) argued that NTK tangent features rotate towards directions predictive of the outputs during training.
@@
Besides the eNTK, other objects involving gradients/curvature have appeared in the mechanistic interpretability literature. As mentioned above, the eNTK (3) is constructed by contracting a pair of Jacobians $J(z)=\nabla_{W}f(z)$ along the parameter direction, forming a kernel in data space (shorthand: $K(z,z^{\prime})=J(z)J(z^{\prime})^{T}=\nabla_{W}f(z)\nabla_{W}f(z^{\prime})^{T}$). The Hessian, the second derivative of the loss function wrt parameters $H=\nabla^{2}_{W}\mathcal{L}$, can be approximated by a pair of Jacobians contracted the other way ($H\approx\sum_{i}J(x_{i})^{T}\,\nabla^{2}_{f}\mathcal{L}(f(x_{i}),y_{i})\,J(x_{i})$), and appears frequently in recent work on singular learning theory (e.g (Bushnaq et al., 2024b)). The Hessian induces a “influence function” kernel on data points $\mathcal{I}(z^{\prime},z)=-\nabla_{W}\mathcal{L}(z^{\prime})^{T}H^{-1}\nabla_{W}\mathcal{L}(z)$ ((Koh & Liang, 2017)) that was recently used to interpret LLMs at scale ((Grosse et al., 2023)). The Jacobian also played a role in the “Local Interaction Basis” algorithm ((Bushnaq et al., 2024a)) which proposed to find interpretable directions in activation space in part by aligning would-be directions with singular directions of the intra-layer Jacobian.
@@
The idea of approximating a neural network by keeping only kernel modes up to an emergent spectral scale can be viewed as a naive application of renormalization ideas to mechanistic interpretability (Greenspan, 2025), with the spectral scale setting a hard UV cutoff.
2 Method: NTK pipeline for feature identification
@@
The main idea that we pursue in this paper is to treat the eNTK as fundamental and interpret its top eigenvectors. One motivation for this idea is that empirically, the spectrum of the eNTK is often very anisotropic. (We’ll see examples of this below.) When true, it means that if the linearized formula (1) is a good approximation to the function learned by a neural network at the end of training, then the same formula where we replace the kernel with its low-rank spectral truncation should in turn be a good approximation, and the top eigenvectors of the kernel should contain most of the information about how the trained model computes its outputs.
Below, for several models, we will eigendecompose the empirical NTK and its variants defined in section 2.1. We define a cliff to be a qualitatively large drop in the ordered eigenvalues  [^2] Alternatively, a more precise definition of a cliff as ending at an eigenvalue $\lambda_{i}$ if the immediate eigengap is more than twice as large as adjacent gaps, $(\lambda_{i+1}-\lambda_{i})/\frac{1}{2}(\lambda_{i}-\lambda_{i-1}+\lambda_{i+2}-\lambda_{i+1})$ >> 22, holds across all examples discussed throughout this paper. However, we believe that the visual evidence we’ll present of spectra with large gaps, often far in excess of a ratio of 2, provides a more compelling, if qualitative, illustration of the phenomenon. , and a NTK feature subspace to be the span of eigenvectors corresponding to the eigenvalues in a cliff.
@@
To compute if a length-$N$ eNTK eigenvector for $N$ the dataset size aligns with a ground-truth feature, we can take the absolute value of its cosine similarity with a length-$N$ “feature vector” whose entries consist of that feature’s activation on each data point. This gives a scalar for each (eigenvector, feature) pair. To visualize how the eigenvectors across a NTK feature subspace align with a selection of features, we will render heatmaps that show this correlation across all pairs.
2.1 Variants of the empirical NTK used in this paper
@@
The eNTK evaluated over all points in a dataset has shape $(N,N,C,C)$ for $N$ the size of the dataset, and $C$ the number of output neurons (classes). To do an eigendecomposition, we must reduce it to two dimensions. Below, we’ll use two collapses of $K$:
@@
Another variant that we’ll use is the layerwise NTK where rather than sum over all model parameters in (3), we sum only over those parameters belonging to a particular layer, before applying either collapse above. This will prove useful for attributing features to a layer.
3 Results for “Toy Models of Superposition”
In our first experiment, we study the alignment of eNTK eigenvectors with ground-truth features in Toy Models of Superposition (TMS) (Elhage et al., 2022).
@@ 3.1 Review of the setup
the dataset consists of NN length-$n$ vectors whose $i$th entry is set to 0 with probability $S$, and otherwise randomly drawn from $U[0,1]$  [^3] By using a fixed-size dataset, we’re actually using the setup of (Henighan et al., 2023). ; and the loss function is MSE loss modified by a per-feature importance hyperparameter $I_{i}$,
@@
The claim to fame of this setup is that it’s a maximally simple toy model of superposition, i.e., the idea that if models store features as linear combinations of within-layer neurons, then a model might learn to represent more features than it has neurons by storing them non-orthogonally in the neural activations. In this experiment, the $n$ input/output directions are the analogs of ground-truth features, and one would like to see if the autoencoder is able to reconstruct them. Upon training this class of model, one finds that when the features are dense (small $S$), the model learns to reconstruct the top $m$ features only, but as one dials up the sparsity, the model is able to recover more and more features, at the cost of encoding them non-orthogonally.
@@
Figure 1: Eigenvalue spectrum of the flattened eNTK (top row) and column-normalized, importance-rescaled flattened eNTK eigenvector-feature indicator heatmaps for TMS models with $n=50$ ground-truth features, importance $I_{i}=0.8^{i}$, (hidden-layer size, sparsity) $m=10,S=0.3$ (left column); $m=10,S=0.9$ (middle column) or $m=40,S=0.9$ (right column); and importance rescaling parameters $\beta=1$ (middle row) and $\beta=0.3$ (bottom row), as described in the main text. The center column is an example of a TMS model in the sparse regime (high superposition), while the left and right columns are two different ways to push the model to the dense regime (low superposition).
@@ 3.2 Results
Spectral structure and cliffs. Across TMS settings (dense vs. sparse regime and different widths $m$), the flattened eNTK exhibits a pronounced spectral cliff structure separating a small set of leading modes from a long tail (Fig. 1, row 1). In all cases that we tested, the leading eigenspace contains two cliffs with the second drop appearing at $k=50$, equal to the number of ground-truth features in the dataset. A first cliff differentiates the number of features that the model is able to reconstruct from those it cannot.  [^4] Quantitatively, in each example the number of entries along the diagonal of $W^{T}W$ exceeding the threshold of 0.75, for $W$ the trained weight matrix, equals the size of one of the two cliffs.
@@
Alignment with ground-truth features. As discussed in section 2, one way to visualize how well a length-$N$ eigenvector for $N$ the dataset size lines up with a ground-truth feature is to compute its inner product with a “feature vector” whose entries consist of that feature’s activation on each data point. Here, we chose instead to work with the flattened eNTK which is a shape $NC$ by $NC$ tensor. At the same time, the output classes $C$ correspond to ground-truth features in TMS. To construct heatmaps displaying how well the flattened eNTK eigenvectors align with ground-truth features in TMS, we therefore first build an expanded data matrix of shape $(NC,C)$ whose entries are zero unless the class indices match, then compute and display the inner product of the expanded data matrix with the top flattened eNTK eigenvectors.
@@
Before rendering the heatmaps, we also apply an importance rescaling to the eNTK, multiplying each Jacobian entry $df_{i}(x)/dW_{\mu}$ in (3) by $I^{\beta i/2}$. This directionally correct for the fact that the derivation of Eq. (1) assumed a standard MSE loss (see Appendix A), but TMS uses an importance-weighted MSE loss, (5). The choice $\beta=1$ would allow us to absorb $I$ into the NTK in eq. (13), allowing the rest of the derivation in Appendix A to proceed with standard MSE, while $\beta=0$ corresponds to turning off the importance rescaling. In practical terms, because the authors of TMS did not model the importance distribution of realistic data, it’s unclear which choice of $\beta$ best matches a realistic data distribution.
@@
With $\beta=1$, we find that the leading eigenvectors align with high-importance features (middle row of Fig. 1). As $\beta$ decreases, lower-importance features become increasingly visible in the heatmaps, until around $\beta=0.3$ (bottom row of Fig. 1), the leading eigendirections of the importance-rescaled eNTK appear to align nearly one-to-one with all of the ground-truth features of TMS in both the sparse and dense regimes. This experiment suggests that the raw kernel also contains information about less-important features, but down-weighs them in line with the objective.
@@
In the sparse (high superposition) regime, the map between top eNTK eigenvectors and ground-truth features appears to split into two lines sorted by importance (center and bottom middle plots in Fig. 1). It could be interesting - and likely is possible given the simplicity of the kernel (see Appendix B for its closed form) - to try to understand why this happens. However, we will leave this for future work.
4 Results for a MLP trained on modular arithmetic